OpenStudy (anonymous):

Tom and Jerry were at a carnival. They each got a ticket for the carousel. Tom got a jumbo pretzel and Jerry got a snow cone. Tom had a certificate for $5 off the cost. Jerry paid the rest, which came to $12.75. Each carousal ticket costs $4 and a snow cone costs $4.50. How much is a jumbo pretzel? A) $7.50 B) $1.50 C) $5.25 D) $3.50

OpenStudy (kaylieanime):

They both bought tickets, so (2*4). One snow cone was bought (4.5) but we don't know the pretzel (x). Tom can take off some with the certificate (-5). The rest Jerry pays (12.75). The equation would be (2*4) + 4.5 + x = 12.75 + 5. Can you simplify it?

OpenStudy (anonymous):

C ?

OpenStudy (kaylieanime):

A good way to check if it's right is plug in it for x. I'm going to simplify it a lot: 12.5 + 5.25 = 17.75 17.75 = 17.75. So Yes! :)